Find efficiency of the cycle

In the Rankine cycle for a steam power plant the turbine entry and exit enthalpies are 2803 kJ/kg and 1800 kJ/kg, respectively. The enthalpies of water at pump entry and exit are 121 kJ/kg and 124 kJ/kg respectively. The specific steam consumption (in kg/kWh) of the cycle is
Answer #1
Turbine work = 2803 - 1800 = 1003 KJ/kg Compressor work = 124-121= 3kJ/kg Net work = 1003-3=1000 kJ/kg SFC = 3600/1000= 3.6 kg/kWh
source: Calculations
answered by: Khaled_bag